OpenStudy (anonymous):

Is this right? Solve the equation. Show your work. 45 = 3b + 69 45- -45 3b 24 3b 24/3 =8 b = 8

rishavraj (rishavraj):

don't u think it should be 3b = -24

OpenStudy (anonymous):

I divided by 3b/3 cancel them out

rishavraj (rishavraj):

0 = 3b + 24 3b = -24

OpenStudy (anonymous):

When i divide by 3 doe i put the 3 on the left or right?

rishavraj (rishavraj):

45 = 3b + 69 45 - 45 = 3b + 69 - 45 0 = 3b + 24 -24 = 3b so b = -8

OpenStudy (anonymous):

How is the 24 negative if it was subtracted by a bigger number 69-45 = 24

rishavraj (rishavraj):

u get 3b + 24 = 0 so 3b + 24 -24 = -24 3b = -24

OpenStudy (anonymous):

Alright the 3 goes on the left. i was putting it on the right that's how i kept getting positive results. Thanks for explaining.